OpenStudy (anonymous):

The specific heat of a certain type of metal is 0.128 J/(g·°C). What is the final temperature if 305 J of heat is added to 86.1 g of this metal initially at 20.0 °C?

OpenStudy (frostbite):

You know how to related the energy given to the heat capacity?

OpenStudy (anonymous):

I do not know how to. What is the equation to do the problem, that's what I cannot remember.

OpenStudy (frostbite):

Right the equation we are going to be using is: \[\Large Q=m \times c \times \Delta T\] Where Q is the heat added, m is the mass, c is the specific heat capacity and \(\Delta T\) is the difference in temperature expressed as \(\Delta T=T_{2}-T_{1}\) .

OpenStudy (frostbite):

And welcome to \(\LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study}}\) by the way.

OpenStudy (frostbite):

Try solve for \(T_{2}\) first.

OpenStudy (frostbite):

Let me help you a bit on the way: \[\Large \Delta T=T _{2}-T _{1}=\frac{ Q }{ m \times c }\]

OpenStudy (anonymous):

Okay, so I am solving for T2 first. T1=20.0, Q= 305 J, m+86.1, and 0.128 J/(g·°C)

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

Oh, and thank you! :)

OpenStudy (frostbite):

That is correct the values you've set. :)

OpenStudy (anonymous):

Yay! So do I need to change 0.128 to just degrees celcius or is it okay as is?

OpenStudy (frostbite):

It should be okay to hold it as celsius as all the units cancel out to celcius in the end.

OpenStudy (frostbite):

But if you come with an answer I'll be more than happy to check it.

OpenStudy (anonymous):

Sorry, my computer died. I got 27.67 for the answer.. and I do not think that that is correct.. :/

OpenStudy (frostbite):

Hey, so did mine. \[\large T _{2}=\frac{ 305 ~ \textrm{J} }{ 86.1 ~ \textrm{g} \times 0.128 \frac{ \textrm{J} }{ \textrm{g} \times ^{\circ} \textrm{C} } }+20.0 ^{\circ} \textrm{C}=27.67 ^{\circ} \textrm{C}+20.0 ^{\circ} \textrm{C}=47.67 ^{\circ} \textrm{C}\] You just forgot to add \(T_{1}\) :)

OpenStudy (frostbite):

I said I would only give a hit so the last equation I gave you should be rearranged to: \[\Large T _{2}=\frac{ Q }{ m \times c }+T_{1}\] But you did it correctly. you calculated how much the temperature would increase from it's current state.